### v2

# Adaptive Contrastive Decoding in Retrieval-Augmented Generation for Handling Noisy Contexts

Abstract
When using large language models (LLMs) in knowledge-intensive tasks, such as open-domain question answering, external context can bridge the gap between external knowledge and the LLMs’ parametric knowledge. Recent research has been developed to amplify contextual knowledge over the parametric knowledge of LLMs with contrastive decoding approaches. While these approaches could yield truthful responses when relevant context is provided, they are prone to vulnerabilities when faced with noisy contexts. We extend the scope of previous studies to encompass noisy contexts and propose adaptive contrastive decoding (ACD) to leverage contextual influence effectively. ACD demonstrates improvements in open-domain question answering tasks compared to baselines, especially in robustness by remaining undistracted by noisy contexts in retrieval-augmented generation.
**footnotetext: Corresponding author.
Adaptive Contrastive Decoding in Retrieval-Augmented Generation for Handling Noisy Contexts
Youna Kim1, Hyuhng Joon Kim1, Cheonbok Park2 3, Choonghyun Park1, Hyunsoo Cho4, Junyeob Kim1, Kang Min Yoo1 2 5, Sang-goo Lee1 6, Taeuk Kim7 * 1Seoul National University, 2NAVER Cloud, 3KAIST AI, 4Ewha Womans University, 5NAVER AI LAB, 6IntelliSys, Korea, 7Hanyang University {anna9812, heyjoonkim, pch330, juny116, sglee}@europa.snu.ac.kr {cbok.park, kangmin.yoo}@navercorp.com, chohyunsoo@ewha.ac.kr kimtaeuk@hanyang.ac.kr
1 Introduction
While large language models (LLMs) (Touvron et al., 2023; Achiam et al., 2023) achieve remarkable performance levels across diverse benchmarks, they sometimes struggle to generalize to knowledge-intensive tasks, such as open-domain question-answering (QA; Chen et al., 2017), and may also fail to capture long-tail knowledge, leading to unfaithful output generation (Mallen et al., 2023; Kandpal et al., 2023). One common approach to address these limitations is fine-tuning the model, but this results in a quadratic rise in computational demands as the size of the LLMs increases exponentially (Longpre et al., 2023). To overcome this, researchers have been investigating strategies to combine non-parametric knowledge with LLMs during response generation without explicit re-training (Asai et al., 2023a). This approach leverages external information from knowledge bases and enhances the capability of the LLMs dynamically, ensuring that the information is both current and accurate.
Figure 1: An illustration of adaptive contrastive decoding (ACD). Entropy ($H$) changes depending on context relevance, affecting the adaptive weight ($\alpha_{\text{ACD}}$). Noisy context leads the model to incorrectly answer "Diede De Groot" when employing regular greedy decoding. ACD applies context-based adjustments, enabling the correct answer, "Sloane Stephens," despite the noise.
Early studies in this field attempt to append query-relevant context to generate more accurate responses. Especially, contrastive decoding (Li et al., 2023; Malkin et al., 2022; Liu et al., 2021) yields significant enhancement in various tasks by amplifying the influence of the given context at decoding step (Shi et al., 2023; Zhao et al., 2024). While such methods work well when context information is correct and faithful, in real-world scenarios, context information is not always correct and may contain some noisy and unfaithful information. For instance, if the retrieval system pulls in irrelevant or contradictory information, it could lead to incorrect responses (Wang et al., 2024; Wu et al., 2024; Yu et al., 2024). This highlights the necessity for a generation model that can gauge the appropriateness of the context by itself, being robust to noise and unfaithful data to ensure the output remains reliable (Yoran et al., 2024).
To assess whether the existing contrastive decoding approaches can be utilized in practice, we extend the setting to situations where the gold-standard context is not guaranteed, specifically in the retrieval-augmented generation (RAG) framework (Yao et al., 2022; Shi et al., 2024; Izacard et al., 2023). In this paper, we demonstrate that existing context-aware contrastive decoding approaches experience performance drops in open-domain question answering, especially when the retrieved context is noisy. To address this issue, we propose adaptive contrastive decoding (ACD), adaptively weighting the contrastive contextual influence on the parametric knowledge, making it suitable for noisy context settings (Figure 1).
Incorporating the distinction between contextual and parametric knowledge, our approach aims to mitigate the dominance of potentially noisy contextual information in model output. We control contrastive contextual influence based on context’s contribution to the LLM’s uncertainty reduction, thereby minimizing its disruptive effect during decoding. Through in-depth experiments with three open-domain QA datasets, we demonstrate the potential of the proposed approach with increased overall performance. Moreover, ACD enhances the performance significantly on the noisy context scenario while minimizing performance degradation on the gold context scenario compared to the baselines.
2 Related Works
Context-Augmented Generation
Approaches for context-augmented generation have been developed to enhance the model’s limited parametric knowledge by providing external knowledge, enabling more factual and contextually accurate responses during inference (Zhou et al., 2023; He et al., 2024). To sufficiently incorporate the information from the context in model generation, contrastive decoding approaches are applied to overwrite the model’s parametric knowledge with external knowledge (Shi et al., 2023; Zhao et al., 2024). These context-aware contrastive decoding methods to generate responses faithful to the given context show effective performance in summarization (See et al., 2017; Narayan et al., 2018), knowledge conflict (Longpre et al., 2022), and question answering with gold-standard contexts.
Robustness in RAG Frameworks
While retrieval-augmented generation enables LLMs to become factual and reliable with the retrieved external knowledge, there are still concerns about incorrectly retrieved irrelevant contexts (Yoran et al., 2024). To address hallucination errors posed by irrelevant contexts, some researchers take an approach to train LLMs that can adaptively retrieve relevant context (Asai et al., 2023b; Wang et al., 2024). Another approach aims to selectively use retrieved contexts after assessing their truthfulness or relevance through context verification with prompting strategies or training untruthful context detectors (Yu et al., 2024; Zhang et al., 2024). These approaches highlight the ongoing efforts to advance the robustness and accuracy of LLMs in multiple directions to manage potentially misleading information.
3 Methodology
3.1 Problem Formulation
At decoding time step $t$, given the input $x$ and preceding sequences $y_{<t}$, a pretrained auto-regressive LLM $\theta$ computes the logit $\mathbf{z}_{t}\in\mathbb{R}^{|V|}$, where $V$ is the vocabulary, for the $t$-th token. In the open-domain QA task, a question $q$ serves as the input $x$, and $\mathbf{z}_{t}$ relies solely on the LLM’s parametric knowledge. When both $q$ and the retrieved context $c$ are provided as $x$, the logit is denoted as $\mathbf{z}_{t}^{c}\in\mathbb{R}^{|V|}$.
3.2 Contrastive Decoding
In cases where context cannot be blindly trusted, directly following the context-augmented distribution can increase the risk of being misled. Thus, we adopt the approach of adding the contextual influence, which contrasts with the LLM’s parametric knowledge, to the parametric distribution $\mathbf{z}_{t}$. With the contrastive decoding objective, $\mathbf{z}_{t}^{c}$ and $\mathbf{z}_{t}$ are ensembled to reflect the influence of external context on the LLM’s parametric knowledge at each decoding step $t$. The probability distribution $P_{\theta}(Y_{t}|x,y_{<t})$ is modified by weighted adjustment based on the difference between $\mathbf{z}_{t}^{c}$ and $\mathbf{z}_{t}$, as represented in the following equation.
$P_{\theta}(Y_{t}\ |\ x,y_{<t})=\text{softmax}(\mathbf{z}_{t}+\alpha\ (\mathbf{z}_{t}^{c}-\mathbf{z}_{t}))$ (1)
The contrastive adjustment enables the LLM to integrate external context $c$ into its prediction, leveraging the weight $\alpha$ to control the impact of $c$ on the final probability distribution.
3.3 Adaptive Weight on Contextual Influence
The degree to which contextual influence is incorporated into $\mathbf{z}_{t}$ needs to be controlled based on the provided context’s informativeness. In practice, however, it is often unknown whether the context is gold or noisy. To address this, we investigate whether the model could adjust accordingly with a simple entropy-based approach.
The LLM’s uncertainty is expressed with the entropy $H(Y_{t})$ of its probability distribution $P_{\theta}(Y_{t}|x,y_{<t})$ (Huang et al., 2023; Kuhn et al., 2023). While $H(Y_{t})$ reflects how much uncertainty the model has based on its parametric knowledge under the given question, $H(Y_{t}^{c})$ is influenced by the external knowledge within the retrieved context $c$. Generally, when the context is added, the entropy decreases (Kendall and Gal, 2017). However, if the context is noisy, irrelevant, or provides no information to answer the given question, it may contribute to increased uncertainty instead.
Intuitively, if the retrieved context provides informative cues for answering the question, then $H(Y_{t}^{c})$ is expected to be lowered compared to $H(Y_{t})$. Conversely, if the context is non-helpful or even confusing the model prediction, $H(Y_{t}^{c})$ in predicting the next token is likely to be higher. This scenario would be particularly evident when the model knows the answer with low $H(Y_{t})$.
Considering the above scenarios, the motivation behind the adaptive weight $\alpha_{ACD}$ is to assign a relatively smaller weight in cases where the context increases uncertainty by being uninformative or confusing for the model in answering the given question. Thus, the value of $\alpha_{ACD}$ is set as the proportion of uncertainty contributed by $H(Y_{t})$ relative to the total uncertainty when considering both $H(Y_{t})$ and $H(Y_{t}^{c})$:
$\alpha_{ACD}=\frac{H(Y_{t})}{H(Y_{t})+H(Y_{t}^{c})}$ (2)
Under the condition where $H(Y_{t})>H(Y_{t}^{c})$, $\alpha_{ACD}$ value approaches to 1, indicating that when the context $c$ is provided, the uncertainty associated with predicting the next token decreases. Conversely, when $H(Y_{t})<H(Y_{t}^{c})$, $\alpha_{ACD}$ value approaches to 0, reflecting minimal influence from $c$. Note that when $H(Y_{t})=H(Y_{t}^{c})$, $\alpha_{ACD}$ becomes 0.5, resulting in an ensemble of two distributions, $\mathbf{z}_{t}$ and $\mathbf{z}_{t}^{c}$, with equal weighting.
With $\alpha_{ACD}$, the vocab $v$ with maximum probability is selected as the next token under the following distribution:
$\hat{P_{\theta}}(Y_{t}\ |\ x,y_{<t})=\text{softmax}(\mathbf{z}_{t}\ +\ \alpha_{ACD}\ (\mathbf{z}_{t}^{c}-\mathbf{z}_{t}))$ (3)
Informed by $\alpha_{ACD}$ and contextual contrast, the adjustment process determines the degree to which the model’s parametric knowledge is superseded, thus optimizing the assimilation of contextual information throughout decoding.
Dataset ($\rightarrow$) TriviaQA NQ PopQA
Model Method ($\downarrow$) All $\text{Subset}_{\text{Gold}}$ $\text{Subset}_{\text{Noisy}}$ All $\text{Subset}_{\text{Gold}}$ $\text{Subset}_{\text{Noisy}}$ All $\text{Subset}_{\text{Gold}}$ $\text{Subset}_{\text{Noisy}}$
Llama2 7B $\text{Reg}_{Cls}$ 59.00 - - 25.48 - - 28.36 - -
$\text{Reg}_{Opn}$ 60.23 87.40 33.50 31.39 61.31 12.40 38.49 81.21 7.77
CAD 49.02 73.69 24.75 25.57 51.61 9.05 33.70 72.18 6.03
MICDF 60.36 85.72 35.39 29.45 56.10 12.54 35.73 74.25 8.03
MICDD 63.23 86.03 40.79 30.36 52.18 16.52 39.01 77.39 11.42
ACD 64.85 88.01 42.06 32.91 56.60 17.88 41.29 82.77 11.46
Llama2 13B $\text{Reg}_{Cls}$ 63.77 - - 30.80 - - 32.70 - -
$\text{Reg}_{Opn}$ 62.81 88.52 37.51 33.35 62.96 14.58 40.03 83.20 8.98
CAD 52.62 76.78 28.85 27.87 55.96 10.05 35.86 76.38 6.71
MICDF 63.53 87.40 40.04 32.63 59.67 15.48 38.16 77.04 10.21
MICDD 66.52 87.68 45.69 34.38 57.32 19.83 41.65 79.27 14.60
ACD 67.37 89.36 45.74 36.12 61.17 20.24 43.35 83.98 14.14
Llama3 8B $\text{Reg}_{Cls}$ 61.67 - - 28.34 - - 32.65 - -
$\text{Reg}_{Opn}$ 61.27 86.94 36.02 33.30 63.10 14.40 39.73 82.95 8.64
CAD 49.70 72.45 27.31 29.17 58.39 10.64 35.86 76.82 6.40
MICDF 61.01 85.40 37.00 27.62 51.89 12.22 37.99 77.12 9.85
MICDD 64.01 86.08 42.28 30.72 53.96 15.98 41.35 79.32 14.04
ACD 66.32 89.20 43.81 35.48 62.03 18.65 43.25 84.48 13.60
Mistral 8B $\text{Reg}_{Cls}$ 63.72 - - 29.64 - - 29.04 - -
$\text{Reg}_{Opn}$ 60.45 86.85 34.48 32.55 64.67 12.18 38.28 81.26 7.36
CAD 44.69 66.89 22.85 24.10 52.25 6.25 33.93 73.95 5.15
MICDF 63.33 88.43 38.62 31.80 61.10 13.22 36.58 76.00 8.23
MICDD 66.97 89.24 45.05 33.24 57.89 17.61 39.87 78.46 12.11
ACD 67.82 90.16 45.83 35.37 62.17 18.38 41.47 82.90 11.68
Table 1: EM accuracy of full data (All) and subsets with gold ($\text{Subset}_{\text{Gold}}$) and noisy contexts ($\text{Subset}_{\text{Noisy}}$). The highest score is in bold, and the second-best is underlined.
4 Experimental Results
4.1 Experimental Settings
Datasets and Models
We conduct experiments on open-domain QA datasets, TriviaQA (Joshi et al., 2017), Natural Questions (NQ; Kwiatkowski et al., 2019), and PopQA (Mallen et al., 2022) with Wikipedia contexts. [^2] Wikipedia dump from Dec. 2018.
We use auto-regressive language models, Llama2 (7B & 13B, Touvron et al., 2023), Llama3 8B, [^3] https://github.com/meta-llama/llama3 and Mistral 7B (Jiang et al., 2023). Utilizing Contriever-msmarco (Izacard et al., 2022) as a retriever, the top-1 retrieved context is appended to each question.
Evaluation Metric
Following Zhao et al. (2024), we use few-shot prompts with 5 examples. We report Exact Match (EM) as an evaluation metric, which verifies whether the generated sequences precisely match one of the candidate answers.
Baselines
As fundamental baselines, regular greedy decoding has been employed in open-book ($\text{Reg}_{Opn}$) and closed-book ($\text{Reg}_{Cls}$) settings. We compare our method against existing context-aware contrastive decoding methods, including Context-Aware Decoding (CAD; Shi et al., 2023) and Multi-Input Contrastive Decoding (MICD; Zhao et al., 2024). MICD uses inputs with and without context, along with an additional input with adversarial context, to generate the output distribution. MICD presents two methods, referred to as MICDF and MICDD, which offer fixed and dynamic $\alpha$, respectively. Similar to our approach, to leverage the burden of hyperparameter search and dependency on fixed $\alpha$, MICDD also determines $\alpha$ dynamically. In $\text{MICD}_{D}$, $\alpha$ is assigned as the maximum token probability with context ($\text{max}P_{wc}$) if $\text{max}P_{wc}$ exceeds the maximum token probability without context ($\text{max}P_{woc}$); otherwise, it is calculated as $1-\text{max}P_{woc}$.
4.2 Main Results
Performance on RAG
As shown in Table 1, ACD outperforms the baselines across all datasets and models within the RAG framework, particularly when considering the full test data (All). When analyzing the performance by dividing the data into two subsets based on whether the retrieved context is gold ($\text{Subset}_{\text{Gold}}$) or not ($\text{Subset}_{\text{Noisy}}$), ACD achieves either the best or second-best performance. $\text{MICD}_{D}$ demonstrates performance comparable to ACD on $\text{Subset}_{\text{Noisy}}$. However, it shows a significant drop on $\text{Subset}_{\text{Gold}}$, indicating a tendency to ignore gold context while handling noisy context. It is notable that both CAD and $\text{MICD}_{F}$ exhibit a significant drop in their performance under noisy conditions.
Performance under Parametric Knowledge
Figure 2: EM accuracy of each method in Llama2-7B. EM of three datasets used are averaged for each subset, Unknown-gold and Known-noisy.
We aim to analyze the model’s performance across various aspects, focusing specifically on its parametric knowledge. We estimate whether the model possesses relevant parametric knowledge for a given question based on its accuracy in a closed-book setting ($\text{Reg}_{Cls}$). We consider two subsets under the following conditions: (1) Known-noisy: the model has parametric knowledge of the given question and noisy context is retrieved. (2) Unknown-gold: the model does not have parametric knowledge of the given question and gold context is retrieved.
From Figure 2, we observe that ACD outperforms the baselines in Known-noisy. Notably, two approaches with adaptively adjusted weight, ACD and $\text{MICD}_{D}$, perform well in Known-noisy, while other baselines show a relative strength in Unknown-gold. However, these baselines also experience significant performance drops in Known-noisy, indicating distraction by noisy context despite correctly answering when only the question is provided. In both cases, ACD demonstrates better performance compared to $\text{MICD}_{D}$, overall showing a tendency towards reliability.
$\alpha$ NQ TriviaQA PopQA
[…]
ACD 65.78 73.37 74.84
[…]
ACD 68.80 72.32 78.90
[…]
ACD 73.27 80.45 80.08
Table 2: AUROC between $\alpha$ used in each method and the noisiness of the retrieved context.
4.3 Analysis
Correlation between Adaptive Weight and Context Noisiness
While other baselines rely on the fixed hyperparameter of weight $\alpha$, ACD and MICDD adjust $\alpha$ during the decoding step. It depends not only on the noisiness of the retrieved context but also on whether the model’s parametric knowledge contains an answer to the given question. To exclude cases that are not directly related to the analysis of how weight is adjusted based on context quality and the model’s parametric knowledge, we use the same subsets, Known-noisy and Unknown-gold.
[…]
Handling Knowledge Conflict
Figure 3: EM accuracy on NQ-swap with contexts replacing the gold answer with a random entity span.
With a knowledge conflict QA dataset, NQ-swap (Longpre et al., 2022), we verify whether the two decoding methods with dynamic weight, ACD and $\text{MICD}_{D}$, can generate context-based responses without considering a conflicting context as a noisy context. The conflicting context in the NQ-swap dataset is constructed by replacing the answer entity span in the original gold context with a random entity of the same type. Figure 3 illustrates that ACD consistently exceeds the performance of $\text{MICD}_{D}$ across all models and achieves results comparable to open-book regular decoding. The results indicate that the ACD’s approach remains effective even in settings where the context is relevant to the question but contradicts the model’s parametric knowledge.
Ablation on $\alpha_{\text{ACD}}$
Figure 4: EM across alpha values ranges from 0.0 to 1.0. The dashed line indicates EM score with $\alpha_{ACD}$.
To assess the impact of $\alpha_{ACD}$ on performance, we fix the value of $\alpha$ within a range $[0,1]$ and examine whether employing ACD is more effective than optimizing a fixed weight. In Figure 4, it can be observed that using a fixed $\alpha$ results in degraded performance compared to ACD. Increasing the alpha value, which enhances the contextual influence on the output distribution, initially leads to a rise in the EM score. However, beyond a certain point, further increasing $\alpha$ results in a decline in the EM score. In scenarios with potential noisy context, a fixed $\alpha$ value may not ensure optimal performance. Therefore, employing an adaptive weight, $\alpha_{ACD}$, to adjust the impact of contextual knowledge based on entropy is crucial for improving overall performance.
5 Conclusion
In this work, we mainly tackle handling noisy contexts in open-domain QA on the RAG framework. Our proposed method, ACD, dynamically adjusts contextual influence during decoding by quantifying the model’s uncertainty that is either reduced or increased by the retrieved context. Our results show that ACD improves performances across various dimensions by considering the LLM’s parametric knowledge and context noisiness. These findings highlight ACD’s potential to enhance the reliability of retrieval-augmented generation.
Limitations
Similar to other contrastive decoding approaches, the inference cost of our approach is higher than the conventional greedy decoding. Specifically, while CAD incurs twice the inference cost and MICD incurs three times the cost, ACD also incurs twice the inference cost of conventional greedy decoding.
Our research is limited the base models and does not encompass chat or instruction-following models trained with reinforcement learning from human feedback (RLHF) or instruction fine-tuning (Ouyang et al., 2022; Chung et al., 2022). These aligned models often generate token distributions that vary significantly based on the presence or absence of contextual instruction or templates. For instance, an instruction-following model might start its generation with "According to the given context …" when context is provided, while directly generating the answer in absence of context. This alignment with the provided instructions poses another challenge to be tackled when the contrastive decoding approach is utilized.
Our current focus is primarily on short-form QA tasks. Expanding to QA tasks with long-form generation will enable a wider range of applications. Under long-form QA tasks, our approach can be further developed to investigate scenarios where the context is only partially relevant to the question.
Acknowledgement
This work was partly supported by SNU-NAVER Hyperscale AI Center and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), No.RS-2020-II201373, Artificial Intelligence Graduate School Program (Hanyang University), NO.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)]
Appendix
Appendix A Implementation Details
A.1 Instructions
The templates we use throughout the experiment are in Table 3 and Table 4. The template used in open-book generation (Table 4) is applied to get context-augmented distribution $\mathbf{z}_{t}^{c}$. Also, to obtain $\mathbf{z}_{t}$, the template in Table 3 is used.
A.2 Datasets
For NQ and TriviaQA, general world knowledge is required to answer the given question. In PopQA, tackling long-tailed information, less popular factual knowledge is asked. For NQ and TriviaQA, few-shot examples are adopted from train data. For PopQA, we randomly sample 5 examples with different relationship types for sample diversity. The number of test data in used is 3,610 for NQ, 11,313 for TriviaQA, and 14,262 for PopQA.
Answer the following questions:
<few-shots>
Question: <question>
Answer:
Table 3: Template used in closed-book generation.
Answer the following questions:
<few-shots>
Context: <context>
Question: <question>
Answer:
Table 4: Template used in open-book generation.
A.3 Baselines
Baselines using regular greedy decoding are evaluated under two different settings. In the closed-book setting, only the question is provided. In the open-book setting, the retrieved context is employed. The same top-1 retrieved context is utilized for every baseline and ACD.
CAD introduces a context-aware contrastive decoding approach that employs a contrastive output distribution to accentuate discrepancies in model predictions with and without context. This method effectively overrides model priors conflicting with provided context, offering significant performance enhancements in tasks requiring resolution of knowledge conflicts. MICD further enhances context grounded generation by integrating contrastive decoding with adversarial irrelevant passages. From a computational time perspective, MICD requires three times more than conventional greedy decoding, while CAD and ACD require twice as much.
MICD proposes two usage directions, referred to as MICDF and MICDD, which offer fixed and dynamic $\alpha$, respectively. MICDD determines $\alpha$ in use by comparing the highest token probabilities with and without given context. Throughout the experiments, fixed value of $\alpha$ is set to the value used in Zhao et al. (2024), 0.5 and 1.0 for CAD and $\text{MICD}_{F}$, respectively.
A.4 Retriever Performance
R@1 R@5 R@10 R@20 R@100
NQ 38.81 65.65 73.91 79.56 88.01
TriviaQA 49.60 71.32 76.72 80.39 85.71
PopQA 41.83 61.54 68.63 74.55 83.95
Table 5: Recall@100 performance for Contriever-msmarco
To assess performance in the RAG framework, the top-1 context from top-100 contexts retrieved by Contriever-msmarco (Izacard et al., 2022) is utilized. Recall@100 is reported for each dataset in Table 5.
A.5 Knowledge Conflict
For the NQ-swap dataset, we utilize the questions and entity-swapped contexts provided in Hong et al. (2024), which includes 3,650 samples. This total excludes 5 few-shot samples and those with contexts presented in a tabular format due to the limited context length. In the case of NQ-swap, each data point has a given context. Since it is a task that does not use a retriever, for MICD, we use the fixed negative context taken from the MICD as an adversarial context. MICD reports that the performance difference between fixed negative and the most distant context is negligible.
Appendix B Results
B.1 Results on Known-noisy and Unknown-gold
For Known-noisy and Unknown-gold, the exact values of EM accuracy on each case are reported in Table 8 and Table 9, respectively.
B.2 AUROC between Adaptive Weight and Context Noisiness
AUROC of ACD and MICDD for three models not reported in Table 2 is reported in Table 10.
Appendix C Additional Analysis
C.1 Upper-bound of Alpha
NQ TriviaQA PopQA
Llama2-7B
$\alpha_{ACD}$ 32.91 64.85 41.29
$\alpha_{oracle}$ 35.35 (+2.44) 65.31 (+0.46) 44.10 (+2.81)
Llama2-13B
$\alpha_{ACD}$ 36.12 67.37 43.35
$\alpha_{oracle}$ 38.75 (+2.63) 68.19 (+0.82) 47.01 (+3.66)
Llama3 8B
$\alpha_{ACD}$ 35.48 66.32 43.25
$\alpha_{oracle}$ 36.98 (+1.50) 66.10 (-0.22) 46.47 (+3.22)
Mistral 7B
$\alpha_{ACD}$ 35.37 67.82 41.47
$\alpha_{oracle}$ 38.37 (+3.00) 67.29 (-0.53) 44.53 (+3.06)
Table 6: EM score comparison between ACD ($\alpha_{ACD}$) and ACD with oracle alpha value ($\alpha_{oracle}$).
In our approach, the parameter $\alpha$ is expected to be close to 1 when the retrieved context contains information that helps answer the given question, and close to 0 otherwise. To evaluate the upper-bound performance of ACD, we assume that we have prior knowledge of whether the context in use is gold or noisy. Under this assumption, we fix the $\alpha$ value to 1.0 if the context is gold and to 0.0 if the context is noisy.
For TriviaQA dataset, the performance of ACD is comparable to $\alpha_{oracle}$, with less than 1 point difference (Table 6). NQ and PopQA show a difference of approximately 2-3 points, indicating that the method for calculating the $\alpha$ weight could be further enhanced in future research.
C.2 Case Study
Sample $\text{Reg}_{Cls}$ $\text{Reg}_{Opn}$ ACD
Case Generation $H(Y_{t})$ Generation $H(Y_{t}^{c})$ Generation $\alpha_{ACD}$
Known-noisy Question: who does the voice Moira Kelly 2.9160 Whoopi Goldberg 5.4562 Moira Kelly 0.3483
of nala in the lion king?
Gold answer: Moira Kelly
Unknown-gold Question: who was the actor that Michael Tucker 6.6748 Michael Moriarty 1.5628 Michael Moriarty 0.8103
played ben stone on law and order?
Gold answer: Michael Moriarty
Table 7: Case study on the value of $\alpha_{ACD}$ for Known-noisy and Unknown-gold cases in Llama2 7B. Each value of entropy without context ($H(Y_{t})$), entropy with context ($H(Y_{t}^{c})$), and $\alpha_{ACD}$ is extracted at the first decoding step ($t=0$).
We conduct the case study on $\alpha_{ACD}$, examining its value in cases of Known-noisy and Unknown-gold. Table 7 shows the generations from Llama2 7B and how the values of entropy from closed-book generation ($\text{Reg}_{Cls}$) and open-book generation ($\text{Reg}_{Opn}$) affect $\alpha_{ACD}$ at the first decoding time step.
In the case of Known-noisy, when the model generates the answer correctly even without the given context, the retrieved noisy context yields relatively higher entropy, resulting in $\alpha_{ACD}$ value of 0.3483. Conversely, in the case of Unknown-gold, the model’s generated answer is incorrect, aligning with a relatively high entropy value of 6.6748. In this scenario, the retrieved gold context guides the model to correctly answer the question, which is reflected in a relatively lower entropy value of 1.5628. Thus, the value of $\alpha_{ACD}$, adjusted with these entropy values, yields a relatively higher weight on the context at 0.8103.
NQ TriviaQA PopQA
Llama2-7B
$\text{Reg}_{Opn}$ 45.13 68.12 33.47
CAD 29.22 48.91 25.97
MICDF 51.07 72.37 36.81
MICDD 72.92 86.33 56.04
ACD 76.72 88.79 54.58
Llama2-13B
$\text{Reg}_{Opn}$ 47.18 69.77 32.53
CAD 32.04 52.48 22.66
MICDF 54.17 75.05 38.55
MICDD 76.31 88.24 59.38
ACD 75.15 88.78 56.11
Llama3-8B
$\text{Reg}_{Opn}$ 46.20 68.50 33.39
CAD 32.91 50.51 23.00
MICDF 43.25 70.67 39.07
MICDD 61.18 83.70 59.80
ACD 64.14 86.59 56.87
Mistral-7B
$\text{Reg}_{Opn}$ 41.04 64.57 31.03
CAD 19.17 42.63 20.80
MICDF 48.12 71.99 36.48
MICDD 69.58 86.84 57.14
ACD 70.62 89.36 53.55
Table 8: EM accuracy of Known-noisy case.
NQ TriviaQA PopQA
Llama2-7B
$\text{Reg}_{Opn}$ 47.78 68.18 74.42
CAD 43.90 62.22 66.12
MICDF 40.47 61.51 64.63
MICDD 29.82 50.43 65.17
ACD 36.03 57.10 73.41
Llama2-13B
$\text{Reg}_{Opn}$ 46.52 65.09 75.04
CAD 45.77 61.07 69.43
MICDF 41.79 62.03 65.85
MICDD 30.72 47.77 64.70
ACD 36.19 53.98 72.38
Llama3-8B
$\text{Reg}_{Opn}$ 48.12 68.47 74.10
CAD 48.00 60.52 70.41
MICDF 38.15 61.40 65.55
MICDD 33.33 48.45 64.81
ACD 41.67 61.24 72.52
Mistral-7B
$\text{Reg}_{Opn}$ 49.57 64.82 73.09
CAD 45.38 56.02 67.81
MICDF 43.28 63.59 66.58
MICDD 32.06 54.97 66.37
ACD 37.73 57.70 73.03
Table 9: EM accuracy of Unknown-gold case.
$\alpha$ NQ TriviaQA PopQA
Llama2 13B
[…]
ACD 69.24 75.31 74.12
[…]
ACD 71.61 73.41 77.92
[…]
ACD 73.07 77.96 80.51
Llama3 8B
[…]
ACD 63.12 57.82 75.00
[…]
ACD 64.08 59.67 75.90
[…]
ACD 67.48 75.45 80.31
Mistral 7B
[…]
ACD 71.27 77.46 74.11
[…]
ACD 76.02 78.20 79.08
[…]
ACD 75.75 84.11 82.07
Table 10: AUROC between $\alpha$ used in each method and the noisiness of the retrieved context. The best AUROC is in bold.